6/11/99 owl shift summary

Jason Hamm (jhamm@physics.arizona.edu)
Fri, 11 Jun 1999 05:04:26 -0700 (MST)

Shift Summary for owl shift, 06/11/99

On Shift: E. Cheu, J. Hamm

Also present: J. LaDue

Hours of beam delivered at all intensities: 7
Hours of 832 data:1
Hours of special runs: 0

Accounting of lost time: Beam studies, DA problems

===================================================================
Runs ended this shift:

13251-13253: BEAM_832
===================================================================
Accesses during shift: none
===================================================================
Tapes brought to FCC this shift: 0
===================================================================
Description of shift:

00:00 Run 13251 in progress; Rick Ford conducting beam studies.

00:15 Writing of accidentals to disk in BEAM_832 configuration disabled
to (hopefully) avoid DART problem on ktev3-will take effect on
next run.

00:50 ktev4 hung-DC2 found to be in DMA state. Run 13251 stopped, fresh
start.

00:57 Heartbeat in DC2, plane 4, stream 2 died. JRL resets DC2-doesn't
work...resets 167 module-doesn't work...reseats board-this works.

1:50 HCC DYC buffer full-fixed by kfrend

2:15 HCC DYC buffer full again-kfrend again

2:35 Rick done with beam studies for the night; run 13522 stopped. JRL
starts tracking down "Missing FFFF word" error from kumquats.
Error isolated to 4Y-9 kumquat. Kumquat replaced, error persists.
JRL concludes error due to bad blossom board. After he replaces
the blossom no more of these errors are seen.

3:00 Persistent pipeline errors-solved only after rebooting all 16
modules.

4:40 Losing stream 5 permit repeatedly-fixed by cycling laser crate
repeater remotely.

5:15 Run 13523 stopped to conduct 832 trigger studies.

5:55 Trigger studies done-run 13524 (BEAM_832) started.